# Convergence Analysis and Acceleration of Fictitious Play for General Mean-Field Games via the Best Response

### v1

@@ 2.2.2 The stability of the best response
Lemma 2.10.
For given $\rho^{i},i=0,1$, let $(\hat{\rho}^{i},\hat{m}^{i})=\operatorname*{argmin}_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho^{i})$, if $H$ is strictly convex and Assumption 2 holds, then
@@
where $D\geq 0$. If in addition, $\hat{\rho}^{0}>0,\hat{\rho}^{1}>0$, then $D=0$ if and only if $\hat{\rho}^{0}=\hat{\rho}^{1}$.
@@
Combining the above gives (19) and by strict convexity of $H$, $D\geq 0$. If $\hat{\rho}^{0}=\hat{\rho}^{1}$ then by definition $D=0$. If $D=0$, then $\nabla\hat{\phi}^{0}=\nabla\hat{\phi}^{1}$. Since $\hat{\rho}^{0}$ and $\hat{\rho}^{1}$ start from the same distribution and are driven by the same vector field $D_{p}H(\nabla\hat{\phi}^{0})=D_{p}H(\nabla\hat{\phi}^{1})$, we have that $\hat{\rho}^{0}=\hat{\rho}^{1}$. ∎
@@
Lemma 2.10 shows that $\rho^{0}-\rho^{1}$ and $\hat{\rho}^{0}-\hat{\rho}^{1}$ are always on the “opposite” side of the “plane” determined by $f(\rho^{0})-f(\rho^{1})$, as illustrated in Figure 1 (left). Moreover, when $\hat{\rho}^{0}$ and $\hat{\rho}^{1}$ are always nonzero, $\hat{\rho}^{0}-\hat{\rho}^{1}$ and $f(\rho^{0})-f(\rho^{1})$ are “orthogonal” if and only if $\hat{\rho}^{0}-\hat{\rho}^{1}$ degenerate. Taking $\rho^{0}=\rho$ be any probability distribution, and $\rho^{1}=\rho^{*}$ be the Nash Equilibrium in Lemma 2.10, Figure 1 (right) shows that $\rho$ and $\hat{\rho}$ are always on the “opposite” sides of $\rho^{*}$ and explains why iteratively conduct best response mapping may not converge to the Nash Equilibrium $\rho^{*}$. The numerical experiment in Section 4.1 verifies that $D\geq 0$ and achieves 0 only when $\hat{\rho}=\rho^{*}$.
While Lemma 2.10 gives a qualitative description, to understand the best response mapping and establish the convergence rate for fictitious play, we need a more careful stability analysis. To be precise, we next show that the best response mapping is stable in three aspects:
  • Lemma 2.11: the gain of the best response at $\rho$ is small, if and only if $\rho$ is close to its best response $\hat{\rho}$.
@@
  • Lemma 2.13: if the costs $f(\rho^{0})$ and $f(\rho^{1})$ are close, then the best replies $\hat{\rho}^{0}$ and $\hat{\rho}^{1}$ are close. Specifically, if $f$ is Lipschitz continuous and $\rho$ is close to the fixed point $\rho^{*}$, then $\rho$ is close to its best response $\hat{\rho}$.
@@
$g(\rho,m)\sim\|\rho-\hat{\rho}\|_{Q_{T}}\sim\|\rho-\rho^{*}\|_{Q_{T}}.$
We first prove that the distance between $\rho$ and its best response $\hat{\rho}$ is equivalent to the gain of the best response at $\rho$. The most important property for this lemma to hold is the constraint being linear. A similar technique is used in [41].
@@
Similarly, if $r$ is directional $L_{r}$-Lipschitz smooth between $(\tilde{\rho},\tilde{m})$ and $(\hat{\rho},\hat{m})$, then
$J(\tilde{\rho},\tilde{m};\rho)-J(\hat{\rho},\hat{m};\rho)\leq\frac{L_{r}}{2}\|\hat{\rho}-\tilde{\rho}\|_{Q_{T}}^{2}.$ (24)
As a consequence, if $r$ is directional $\lambda_{r}$-strongly convex and directional $L_{r}$-Lipschitz smooth between $(\rho,m)$ and its best response $(\hat{\rho},\hat{m})$, then the distance between $\rho$ and its best response is equivalent to the gain of the best response at $\rho$, i.e.
$\frac{\lambda_{r}}{2}\|\hat{\rho}-\rho\|^{2}_{Q_{T}}\leq g(\rho,m)\leq\frac{L_{r}}{2}\|\hat{\rho}-\rho\|^{2}_{Q_{T}}.$ (25)
@@
$J(\tilde{\rho},\tilde{m};\rho)-J(\hat{\rho},\hat{m};\rho)=r(\tilde{\rho},\tilde{m})-r(\hat{\rho},\hat{m})+\langle f(\rho),\tilde{\rho}-\hat{\rho}\rangle_{Q_{T}}+\langle f_{T}(\rho_{T}),\tilde{\rho}_{T}-\hat{\rho}_{T}\rangle_{\Omega}$ (26)
The directional strong convexity of $r$ gives that
@@
And the directional Lipschitz smoothness gives that
$r(\tilde{\rho},\tilde{m})-r(\hat{\rho},\hat{m})\leq\left\langle L(x,\frac{\hat{m}}{\hat{\rho}})-\frac{\hat{m}}{\hat{\rho}}D_{v}L(x,\frac{\hat{m}}{\hat{\rho}}),\tilde{\rho}-\hat{\rho}\right\rangle_{Q_{T}}+\left\langle D_{v}L(x,\frac{\hat{m}}{\hat{\rho}}),\tilde{m}-\hat{m}\right\rangle_{Q_{T}}+\frac{L_{r}}{2}\|\tilde{\rho}-\hat{\rho}\|_{Q_{T}}^{2}.$ (28)
@@
$A=\langle-\partial_{t}\hat{\phi}-\nu\Delta\hat{\phi},\tilde{\rho}-\hat{\rho}\rangle_{Q_{T}}+\langle-\nabla\hat{\phi},\tilde{m}-\hat{m}\rangle_{Q_{T}}+\langle f_{T}(\rho_{T}),\tilde{\rho}_{T}-\hat{\rho}_{T}\rangle_{\Omega}$ (32)
@@
Therefore (23) and (24) are true. Taking $(\tilde{\rho},\tilde{m})=(\rho,m)$ and with the definition of gain of the best response in (15), we conclude (25).
The previous lemma shows that if the gain of the best response at $\rho$ is small, then $\rho$ is close to its best response $\hat{\rho}$. Moreover, the next lemma shows that if the difference between $\rho$ and its best response $\hat{\rho}$ is small, then $\rho$ is close to the Nash Equilibrium $\rho^{*}$.
@@
For given $(\rho,m)\in C$, let its best response be $(\hat{\rho},\hat{m})$ and let $(\rho^{*},m^{*},\phi^{*})$ be a Nash Equilibrium. Under Assumptions 1-2, if $f$ is $\lambda_{f}$-strongly monotone, then
@@
Since $(\rho^{*},m^{*},\phi^{*})$ is a Nash Equilibrium, by Lemma 2.7, we know that $(\rho^{*},m^{*})=\operatorname*{argmin}_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho^{*})$. Applying Lemma 2.13 then gives
@@
Let $(\hat{\rho}^{i},\hat{m}^{i})$ be the best response to given $(\rho^{i},m^{i}),i=0,1$. Under the Assumption 2, if $r$ is $\lambda_{r}$-strongly convex between $(\hat{\rho}^{0},\hat{m}^{0})$ and $(\hat{\rho}^{1},\hat{m}^{1})$, then
@@
Denote $(\hat{\rho}^{\epsilon},\hat{m}^{\epsilon}):=(1-\epsilon)(\hat{\rho}^{0},\hat{m}^{0})+\epsilon(\hat{\rho}^{1},\hat{m}^{1})$. Then the optimality of $(\hat{\rho}^{0},\hat{m}^{0})$ gives
$r(\hat{\rho}^{0},\hat{m}^{0})+\langle f(\rho^{0}),\hat{\rho}^{0}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}^{0}_{T}\rangle_{\Omega}\leq r(\hat{\rho}^{\epsilon},\hat{m}^{\epsilon})+\langle f(\rho^{0}),\hat{\rho}^{\epsilon}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}^{\epsilon}_{T}\rangle_{\Omega},$ (41)
@@
$r(\hat{\rho}^{0},\hat{m}^{0})\leq r(\hat{\rho}^{\epsilon},\hat{m}^{\epsilon})+\epsilon\langle f(\rho^{0}),\hat{\rho}^{1}-\hat{\rho}^{0}\rangle_{Q_{T}}+\epsilon\langle f_{T},\hat{\rho}^{1}_{T}-\hat{\rho}^{0}_{T}\rangle_{\Omega}.$ (42)
@@
$r(\hat{\rho}^{\epsilon},\hat{m}^{\epsilon})+\frac{\epsilon(1-\epsilon)\lambda_{r}}{2}\|\hat{\rho}^{0}-\hat{\rho}^{1}\|^{2}_{Q_{T}}\leq(1-\epsilon)r(\hat{\rho}^{0},\hat{m}^{0})+\epsilon r(\hat{\rho}^{1},\hat{m}^{1}).$ (43)
@@
$\frac{(1-\epsilon)\lambda_{r}}{2}\|\hat{\rho}^{0}-\hat{\rho}^{1}\|^{2}_{Q_{T}}\leq-r(\hat{\rho}^{0},\hat{m}^{0})+r(\hat{\rho}^{1},\hat{m}^{1})+\langle f(\rho^{0}),\hat{\rho}^{1}-\hat{\rho}^{0}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}^{1}_{T}-\hat{\rho}^{0}_{T}\rangle_{\Omega}.$ (44)
Then we let $\epsilon\to 0$ and obtain
$\frac{\lambda_{r}}{2}\|\hat{\rho}^{0}-\hat{\rho}^{1}\|^{2}_{Q_{T}}\leq-r(\hat{\rho}^{0},\hat{m}^{0})+r(\hat{\rho}^{1},\hat{m}^{1})+\langle f(\rho^{0}),\hat{\rho}^{1}-\hat{\rho}^{0}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}^{1}_{T}-\hat{\rho}^{0}_{T}\rangle_{\Omega}.$ (45)
@@
$\frac{\lambda_{r}}{2}\|\hat{\rho}^{0}-\hat{\rho}^{1}\|^{2}_{Q_{T}}\leq-r(\hat{\rho}^{1},\hat{m}^{1})+r(\hat{\rho}^{0},\hat{m}^{0})+\langle f(\rho^{1}),\hat{\rho}^{0}-\hat{\rho}^{1}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}^{0}_{T}-\hat{\rho}^{1}_{T}\rangle_{\Omega}.$ (46)
@@
For given $(\rho,m)\in C$, let its best response be $(\hat{\rho},\hat{m})$ and let $(\rho^{*},m^{*},\phi^{*})$ be a Nash Equilibrium. Under the Assumption 1-2, if $r$ is $\lambda_{r}$-strongly convex between $({\rho},{m})$ and $({\rho}^{*},{m}^{*})$, then
@@
By Assumption 1, $\|f(\rho^{1})-f(\rho^{0})\|_{Q_{T}}\leq L_{f}\|\rho^{1}-\rho^{0}\|_{Q_{T}}$. Taking $\rho^{0}=\rho$ and $\rho^{1}=\rho^{*}$ in Lemma 2.13 and by triangle inequality concludes the proof. ∎
We close this section with remarks on the assumptions.
Remark 2.15 (Independent terminal cost).
The assumption that $f_{T}(x,\rho_{T})$ being independent of $\rho_{T}$ is quite common in MFG literature, see [14, 17, 2, 7, 41, 34] for example. When this assumption holds, the terminal condition of the HJB equation is independent of $\rho_{T}$. But as long as $f(x,\rho)$ depends on $\rho$, the MFG system (2) is still coupled and the problem is not oversimplified.
Remark 2.16 (Strong monotonicity of $f$).
$f$ being strongly monotone is a common assumption in game theory and is related to the price of anarchy [24].
Remark 2.17 (Directional strong convexity of $r$).
For given $(\rho^{0},m^{0}),(\rho^{1},m^{1})\in C$, we can assume that $\frac{m^{0}}{\rho^{0}}\neq\frac{m^{1}}{\rho^{1}}$, otherwise $\rho_{0}=\rho_{1}$ since they are induced by the same velocity field. If in addition, $\rho^{0}>0,\rho^{1}>0$, then it is possible to find $\lambda_{r}>0$ such that $r$ is directional $\lambda_{r}$-strongly convex between $(\rho^{0},m^{0})$ and $(\rho^{1},m^{1})$. To see this, we take $H(x,p)=\frac{1}{2}|p|^{2}$ and $r(\rho,m)=\int_{Q_{T}}\frac{|m|^{2}}{2\rho}$ as an example. With the above assumptions, the second-order derivative of $y$ defined in Definition 2.8 is
$y{{}^{\prime\prime}}(\epsilon)=\int_{Q_{T}}\frac{(\rho^{0}\rho^{1})^{2}}{(\rho^{\epsilon})^{3}}\left|\frac{m^{0}}{\rho^{0}}-\frac{m^{1}}{\rho^{1}}\right|^{2}\geq\int_{Q_{T}}\min\left\{\frac{(\rho^{0})^{2}}{\rho^{1}},\frac{(\rho^{1})^{2}}{\rho^{0}}\right\}\left|\frac{m^{0}}{\rho^{0}}-\frac{m^{1}}{\rho^{1}}\right|^{2}>0.$
Since $y^{\prime\prime}(\epsilon)$ has a positive lower-bound on $[0,1]$, then there exists $\lambda_{r}>0$ such that $y^{\prime\prime}(\epsilon)\geq\lambda_{r}\|\rho^{0}-\rho^{1}\|_{Q_{T}}^{2}$, i.e. $r$ is directional $\lambda_{r}$-strongly convex between $(\rho^{0},m^{0})$ and $(\rho^{1},m^{1})$. The value of $\lambda_{r}$ depends on the choices of two endpoints. But notice that $y^{\prime\prime}(\epsilon)\sim\|v^{0}-v^{1}\|_{Q_{T}}^{2}$, it is possible that $\lambda_{r}=\mathcal{O}(1)$ with respect to $\|\rho^{0}-\rho^{1}\|_{Q_{T}}^{2}$.
Remark 2.18 (Directional strong convexity of $r$ from modified cost).
We can also obtain the directional strong convexity of $r$ if $r(\rho,m)-\frac{\lambda_{r}}{2}\|\rho\|^{2}_{Q_{T}}$ is convex. However, such $\lambda_{r}>0$ may not exist for the original definition of $r$. Nevertheless, we can adapt the definition of $r$ without changing the best response mapping too much if $f$ is $2\lambda_{f}$-strongly monotone. With strong monotonicity of $f$, we let $\lambda_{r}=\lambda_{f}$ and the modified cost be
$\tilde{r}(\tilde{\rho},\tilde{m}):=r(\tilde{\rho},\tilde{m})+\frac{\lambda_{r}}{2}\|\tilde{\rho}\|_{Q_{T}}^{2},\quad\tilde{f}(x,\rho):=f(x,\rho)-\lambda_{r}\rho(x),\quad\tilde{J}(\tilde{\rho},\tilde{m};\rho):=\tilde{r}(\tilde{\rho},\tilde{m})+\langle\tilde{f}(\rho),\tilde{\rho}\rangle_{Q_{T}}+\langle f_{T},\tilde{\rho}_{T}\rangle_{\Omega}.$
It is easy to see that the new costs $\tilde{r}$ and $\tilde{f}$ satisfy that $\tilde{r}-\frac{\lambda_{r}}{2}\|\rho\|^{2}_{Q_{T}}$ is convex and $\tilde{f}$ is $\lambda_{f}$-monotone. Therefore the analysis in this section holds for
With mild assumptions, $(\hat{\rho},\hat{m})=\operatorname*{argmin}_{(\tilde{\rho},\tilde{m})\in C}\tilde{J}(\tilde{\rho},\tilde{m};\rho)$ if $\hat{m}=-\hat{\rho}D_{p}H(\cdot,\nabla\hat{\phi})$ and $(\hat{\rho},\hat{\phi})$ solves
$\hat{\phi}(x,T)=f_{T}(x,\rho(T)),\quad\hat{\rho}(x,0)=\rho_{0}(x),$ $x\in\Omega.$
In other words, when $\rho$ is close to the Nash Equilibrium, the difference $\hat{\rho}-\rho$ is small, and the best response is approximately solving (49), where the assumptions hold. Here the value of $\lambda_{r}$ does not depend on $\rho,m$.
2.2.3 Non-potential MFGs
In this section, we establish the convergence of scheme (16)-(17) and thus fictitious play (Algorithm 1) for non-potential MFGs by adopting the gain of the best response $g(\rho,m)$ as the Lyapunov function.

### v2

@@ 3.3 The stability of the best response
Proposition 3.11.
For given $\rho^{i},i=0,1$, let $(\hat{\rho}^{i},\hat{m}^{i})=\operatorname*{argmin}_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho^{i})$, if Assumption 3.6 holds, then
@@
where $D\geq 0$ and $D=0$ if and only if $\hat{\rho}^{0}=\hat{\rho}^{1}$.
@@
Combining the above gives (33) and by strict convexity of $H$, $D\geq 0$. If $\hat{\rho}^{0}=\hat{\rho}^{1}$ then by definition $D=0$. If $D=0$, then by positivity of $\hat{\rho}^{0}$ and $\hat{\rho}^{1}$, we have $\nabla\hat{\phi}^{0}=\nabla\hat{\phi}^{1}$. Since $\hat{\rho}^{0}$ and $\hat{\rho}^{1}$ start from the same distribution and are driven by the same vector field $D_{p}H(\nabla\hat{\phi}^{0})=D_{p}H(\nabla\hat{\phi}^{1})$, we have that $\hat{\rho}^{0}=\hat{\rho}^{1}$. ∎
Proposition 3.11 shows that when $f$ is monotone, the vectors $\rho^{0}-\rho^{1}$ and $\hat{\rho}^{0}-\hat{\rho}^{1}$ always lie on opposite sides of the “plane” defined by $f(\rho^{0})-f(\rho^{1})$, as illustrated in Figure 1 (left). Furthermore, since $\hat{\rho}^{0}>0$ and $\hat{\rho}^{1}>0$, the vectors $\hat{\rho}^{0}-\hat{\rho}^{1}$ and $f(\rho^{0})-f(\rho^{1})$ are “orthogonal” if and only if $\hat{\rho}^{0}-\hat{\rho}^{1}$ degenerates. The experiment in Section 5.2 verifies that the quantity $D$ remains nonnegative and reaches zero only when $\hat{\rho}=\rho^{*}$.
Specializing Proposition 3.11 to the case where $\rho^{0}=\rho$ is any density evolution and $\rho^{1}=\rho^{*}$ is the Nash equilibrium, Figure 1 (right) illustrates that $\rho$ and $\hat{\rho}$ always lie on opposite sides of $\rho^{*}$. This geometric interpretation helps explain why repeated application of the best response mapping may fail to converge to the Nash equilibrium $\rho^{*}$. The numerical experiment in Section 5.1 confirms this behavior, showing that fixed-point iteration can oscillate between two states and fail to converge.
@@
While Proposition 3.11 gives a qualitative description, to understand the best response mapping and establish the convergence rate for fictitious play, we need a more careful stability analysis. To be precise, we next show that the best response mapping is stable in three aspects:
  • Lemma 3.12: the gain of the best response at $\rho$ is small, only if $\rho$ is close to its best response $\hat{\rho}$.
@@
  • Lemma 3.14: if the costs $f(\rho^{0})$ and $f(\rho^{1})$ are close, then the best responses $\hat{\rho}^{0}$ and $\hat{\rho}^{1}$ are close. Specifically, if $f$ is Lipschitz continuous (Assumption 3.5) and $\rho$ is close to the fixed point $\rho^{*}$, then $\rho$ is close to its best response $\hat{\rho}$.
@@
$c_{(\rho,m)}g(\rho,m)\geq\|\rho-\hat{\rho}\|_{Q_{T}}\sim\|\rho-\rho^{*}\|_{Q_{T}},$
for some $c_{(\rho,m)}\in\mathbb{R}$.
We first prove that the distance between $\rho$ and its best response $\hat{\rho}$ is bounded by the gain of the best response at $\rho$. The key property for this lemma to hold is the constraint being linear. A similar technique is used in Lavigne2023gcg-ficplay .
@@
As a consequence, if $R$ is directional $\lambda_{R}$-strongly convex between $(\rho,m)$ and its best response $(\hat{\rho},\hat{m})$, then the distance between $\rho$ and its best response is equivalent to the gain of the best response at $\rho$, i.e.
$\frac{\lambda_{R}}{2}\|\hat{\rho}-\rho\|^{2}_{Q_{T}}\leq g(\rho,m).$ (38)
@@
$J(\tilde{\rho},\tilde{m};\rho)-J(\hat{\rho},\hat{m};\rho)=R(\tilde{\rho},\tilde{m})-R(\hat{\rho},\hat{m})+\langle f(\rho),\tilde{\rho}-\hat{\rho}\rangle_{Q_{T}}+\langle f_{T}(\rho_{T}),\tilde{\rho}_{T}-\hat{\rho}_{T}\rangle_{\mathbb{T}^{d}}$ (39)
By Lemma 2.4, $\hat{\rho}>0$. Therefore, the directional strong convexity of $R$ gives that
@@
$A=\langle-\partial_{t}\hat{\phi}-\nu\Delta\hat{\phi},\tilde{\rho}-\hat{\rho}\rangle_{Q_{T}}+\langle-\nabla\hat{\phi},\tilde{m}-\hat{m}\rangle_{Q_{T}}+\langle f_{T}(\rho_{T}),\tilde{\rho}_{T}-\hat{\rho}_{T}\rangle_{\mathbb{T}^{d}}$ (43)
@@
Therefore (37) is true. Taking $(\tilde{\rho},\tilde{m})=(\rho,m)$ and with the definition of gain of the best response in (23), we conclude (38).
The next lemma shows that if the difference between $\rho$ and its best response $\hat{\rho}$ is small, then $\rho$ is close to the Nash Equilibrium $\rho^{*}$.
@@
For given $(\rho,m)\in C$, let its best response be $(\hat{\rho},\hat{m})$ and let $(\rho^{*},m^{*},\phi^{*})$ be a Nash Equilibrium. Under Assumptions 3.5-3.6, if in addition $f$ is $\lambda_{f}$-strongly monotone, then
@@
Since $(\rho^{*},m^{*},\phi^{*})$ is a Nash Equilibrium, by Lemma 3.4, we know that $(\rho^{*},m^{*})=\operatorname*{argmin}_{(\tilde{\rho},\tilde{m})\in C}J(\tilde{\rho},\tilde{m};\rho^{*})$. Applying Proposition 3.11 then gives
@@
Let $(\hat{\rho}^{i},\hat{m}^{i})$ be the best response to given $(\rho^{i},m^{i}),i=0,1$. Under the Assumption 3.6, if $R$ is $\lambda_{R}$-strongly convex between $(\hat{\rho}^{0},\hat{m}^{0})$ and $(\hat{\rho}^{1},\hat{m}^{1})$, then
@@
Denote $(\hat{\rho}^{\theta},\hat{m}^{\theta}):=(1-\theta)(\hat{\rho}^{0},\hat{m}^{0})+\theta(\hat{\rho}^{1},\hat{m}^{1})$. Then the optimality of $(\hat{\rho}^{0},\hat{m}^{0})$ gives
$R(\hat{\rho}^{0},\hat{m}^{0})+\langle f(\rho^{0}),\hat{\rho}^{0}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}^{0}_{T}\rangle_{\mathbb{T}^{d}}\leq R(\hat{\rho}^{\theta},\hat{m}^{\theta})+\langle f(\rho^{0}),\hat{\rho}^{\theta}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}^{\theta}_{T}\rangle_{\mathbb{T}^{d}},$ (52)
@@
$R(\hat{\rho}^{0},\hat{m}^{0})\leq R(\hat{\rho}^{\theta},\hat{m}^{\theta})+\theta\langle f(\rho^{0}),\hat{\rho}^{1}-\hat{\rho}^{0}\rangle_{Q_{T}}+\theta\langle f_{T},\hat{\rho}^{1}_{T}-\hat{\rho}^{0}_{T}\rangle_{\mathbb{T}^{d}}.$ (53)
@@
$R(\hat{\rho}^{\theta},\hat{m}^{\theta})+\frac{\theta(1-\theta)\lambda_{R}}{2}\|\hat{\rho}^{0}-\hat{\rho}^{1}\|^{2}_{Q_{T}}\leq(1-\theta)R(\hat{\rho}^{0},\hat{m}^{0})+\theta R(\hat{\rho}^{1},\hat{m}^{1}).$ (54)
@@
$\frac{(1-\theta)\lambda_{R}}{2}\|\hat{\rho}^{0}-\hat{\rho}^{1}\|^{2}_{Q_{T}}\leq-R(\hat{\rho}^{0},\hat{m}^{0})+R(\hat{\rho}^{1},\hat{m}^{1})+\langle f(\rho^{0}),\hat{\rho}^{1}-\hat{\rho}^{0}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}^{1}_{T}-\hat{\rho}^{0}_{T}\rangle_{\mathbb{T}^{d}}.$ (55)
Then we let $\theta\to 0$ and obtain
$\frac{\lambda_{R}}{2}\|\hat{\rho}^{0}-\hat{\rho}^{1}\|^{2}_{Q_{T}}\leq-R(\hat{\rho}^{0},\hat{m}^{0})+R(\hat{\rho}^{1},\hat{m}^{1})+\langle f(\rho^{0}),\hat{\rho}^{1}-\hat{\rho}^{0}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}^{1}_{T}-\hat{\rho}^{0}_{T}\rangle_{\mathbb{T}^{d}}.$ (56)
@@
$\frac{\lambda_{R}}{2}\|\hat{\rho}^{0}-\hat{\rho}^{1}\|^{2}_{Q_{T}}\leq-r(\hat{\rho}^{1},\hat{m}^{1})+r(\hat{\rho}^{0},\hat{m}^{0})+\langle f(\rho^{1}),\hat{\rho}^{0}-\hat{\rho}^{1}\rangle_{Q_{T}}+\langle f_{T},\hat{\rho}^{0}_{T}-\hat{\rho}^{1}_{T}\rangle_{\mathbb{T}^{d}}.$ (57)
@@
For given $(\rho,m)\in C$, let its best response be $(\hat{\rho},\hat{m})$ and let $(\rho^{*},m^{*},\phi^{*})$ be a Nash Equilibrium. Under the Assumption 3.5-3.6, if $R$ is $\lambda_{R}$-strongly convex between $({\rho},{m})$ and $({\rho}^{*},{m}^{*})$, then
@@
By Assumption 3.5, $\|f(\rho^{1})-f(\rho^{0})\|_{Q_{T}}\leq L_{f}\|\rho^{1}-\rho^{0}\|_{Q_{T}}$. Taking $\rho^{0}=\rho$ and $\rho^{1}=\rho^{*}$ in Lemma 3.14 and by triangle inequality concludes the proof. ∎
3.4 Non-potential MFGs
In this section, we adopt the gain of the best response $g(\rho,m)$ as the Lyapunov function to establish the convergence of scheme (24)-(25) and thus fictitious play (Algorithm 1) for non-potential MFGs.